# Limit Theorems and Quantitative Statistical Stability for the Equilibrium States of Piecewise Partially Hyperbolic Maps

### v1

@@ 1. Introduction
  • Generalized Potentials and Dynamics: We extend the statistical analysis to a significantly broader class of potentials $\overline{\mathcal{P}}$ and dynamical systems than those treated in [3]. Notably, our results encompass potentials with explicit dependence on the fiber direction, thereby eliminating the restrictive requirement of fiber-constancy;
@@ 2. Settings and Examples
Let $\mathcal{PH}$ be the set of Hölder potentials which satisfies the following condition (f3) which is open relatively to the Hölder norm.
@@
    There exists a sufficiently small $\epsilon_{\varphi}>0$ s.t.
    $\sup\varphi-\inf\varphi<\epsilon_{\varphi};$ (2)
@@
Where we suppose the constants $\epsilon_{\varphi}$ and $L$ satisfy the condition
$\exp{\epsilon_{\rho}}\cdot\left(\dfrac{(\deg(f)-q)\sigma^{-\alpha}+qL^{\alpha}[1+(L-1)^{\alpha}]}{\deg(f)}\right)<1.$ (4)
@@ 2.2. Spectral gap for $\mathcal{L}_{\varphi}$
      For every $\varphi\in\mathcal{PH}$, there exists an equilibrium state $m(=m(\varphi))$, such that $m=h\nu$, where $0<\inf h\leq\sup h<+\infty$, (see (P1.2)) $h,\frac{1}{h}\in B_{s}$ and $\mathcal{L}_{\varphi}(h)=\lambda h$, $\lambda>0$. Moreover, $\nu$ is a conformal measure with Jacobian $\lambda e^{-\varphi}$;
@@
      There exist normed spaces of real valued functions (not necessarily Banach), closed under the standard multiplication, $(B_{s},|\cdot|_{s})\subset(B_{w},|\cdot|_{w})$ satisfying $|\cdot|_{w}\leq|\cdot|_{s}$ such that for every $\varphi\in\mathcal{PH}$ the spectrum of the Ruelle-Perron-Frobenius of $f:M\longrightarrow M$, $\mathcal{L}_{\varphi}:B_{s}\longrightarrow B_{s}$, defined by
@@
Moreover, the following theorems 2.2 and 2.3 hold, where the proof can be found in [4].
@@
Note that, the probability $m$ is an equilibrium state for a potential $\varphi\in\mathcal{PH}$. Thus $\mu_{0}$ depends on the function $\Phi=\varphi\circ\pi_{1}\in\mathcal{P}$ (see equation (24)).
2.4. Examples.
In this section we expose some examples satisfying all results of this work.
@@
Now we present a general idea to generate examples by perturbing the identity or an expanding map close to the identity.
Let $f_{0}:\mathbb{T}^{d}\longrightarrow\mathbb{T}^{d}$ be an expanding map. Choose a covering $\mathcal{P}$ and an atom $P_{1}\in\mathcal{P}$ that contains a periodic point (possibly a fixed point) $p$. Next, consider a perturbation $f$, of $f_{0}$, within $P_{1}$ using a pitchfork bifurcation in such a way that $p$ becomes a saddle point for $f$. Consequently, $f$ coincides with $f_{0}$ in $P_{1}^{c}$, where we have uniform expansion. The perturbation can be designed to satisfy condition (f1). This ensures that $f$ is never overly contracting in $P_{1}$, and it remains topologically mixing. However, it is important to note that a small perturbation with these properties may not always exist. If such a perturbation does exist, then condition (f3) can be satisfied. In this case, $m$ is absolutely continuous with respect to the Lebesgue measure, which is an expanding conformal and positive measure on open sets. Consequently, there can be no periodic attractors.
@@
In the previous example, assume that $f_{0}$ is diagonalizable, with eigenvalues $1<1+a<\lambda$, associated with $e_{1},e_{2}$, respectively, where $x_{0}$ is a fixed point. Fix $a,\epsilon>0$ such that $\log(\frac{1+a}{1-a})<\epsilon$ and
$\operatorname{e}^{\epsilon}\left(\frac{(\deg(f_{0})-1)(1+a)^{-\zeta}+(1/(1-a))^{\zeta}[1+(a/(1-a))^{\zeta}]}{\deg(f_{0})}\right)<1.$
Note that any smaller $a>0$ will still satisfy these equations.
Let $\mathcal{U}$ be a finite covering of $M$ by open domains of injectivity for $f_{0}$. Redefining sets in $\mathcal{U}$, we may assume $x_{0}=(m_{0},n_{0})$ belongs to exactly one such domain $U$. Let $r>0$ be small enough that $B_{2r}(x_{0})\subset U$. Define $\rho=\eta_{r}\ast g$, where $\eta_{r}(z)=(1/r^{2})\eta(z/r)$, $\eta$ denotes the standard mollifier, and
$g(m,n)=\begin{cases}\lambda(1-a),&\text{if }(m,n)\in B_{r}(x_{0});\\ \lambda(1+a),&\text{otherwise.}\end{cases}$
@@
$f(m,n)=(m_{0}+\lambda(m-m_{0}),n_{0}+(\rho(m,n)/\lambda)(n-n_{0})).$
Then $x_{0}$ is a saddle point of $f$ and the desired conditions are satisfied for $\mathcal{A}=B_{2r}(x_{0})$, $L_{1}=1/(1-a)$ and $\sigma=1+2a$. The only non-trivial condition is (f3). To show it, note that
@@
where $S=\{z\in\mathbb{R}^{2}:x-z\in B_{r}(x_{0}),y-z\notin B_{r}(x_{0})\}$ and $S^{\prime}=\{z\in\mathbb{R}^{2}:y-z\in B_{r}(x_{0}),x-z\notin B_{r}(x_{0})\}$. Take $x,y\in\mathbb{R}^{2}$ and write |xy|=qr|x-y|=qr, $A_{q}=\{z\in\mathbb{R}^{2}:1-q<|z|<1\}$. We have
@@
Since $N=\sup_{q>0}\eta(A_{q})/q^{\zeta}<+\infty$, we can take $a$ so small that $2aN/(1-a)<\epsilon$, therefore $H_{\zeta}(\rho)<\epsilon\inf\rho$.
Example 2.4.
(Manneville-Poumeau map) If $\alpha\in(0,1)$, let $f_{\alpha}:[0,1]\longrightarrow[0,1]$ be the $C^{1+\alpha}$ local diffeomorphism given by
$f_{\alpha}(x)=\left\{\begin{array}[]{ll}x\left(1+2^{\alpha}x^{\alpha}\right)&\text{ if }0\leqslant x\leqslant\frac{1}{2},\\ 2x-1&\text{ if }\frac{1}{2}<x\leqslant 1.\end{array}\right$.
Note that (f1) and (f2) are satisfied. Moreover, computing the two inverse branches of $f_{\alpha}$, it can be shown that $L=1$. For (f3) we consider the following family of potentials $\{\varphi_{\alpha,t}\}_{t\in(-t_{0},t_{0})}$, given by $\varphi_{\alpha,t}=-t\log|Df_{\alpha}|$, for sufficient small $t_{0}$. This is a family of $C^{\alpha}$ potentials such that $\{\varphi_{\alpha,t}\}_{t\in(-t_{0},t_{0})}\subset\mathcal{PH}$ for each $\alpha\in(0,1)$, as we see ahead. Indeed,
$|\varphi_{\alpha,t}(x)-\varphi_{\alpha,t}(y)|$ $=$ $|t\log|Df_{\alpha}(x)|-t\log|Df_{\alpha}(y)||$
$=$ $|t|\log\dfrac{|Df_{\alpha}(x)|}{|Df_{\alpha}(y)|}$
$\leq$ $|t|\log(2+\alpha).$
Example 2.5.
(Discontinuous Maps) Let $F=(f_{\alpha},G)$ be the measurable map, where $f_{\alpha}$ is from the previous example 2.4. Consider the real numbers $\alpha_{1}$ and $\alpha_{2}$ s.t $0\leq\alpha_{1}<\alpha_{2}<1$. Let $G:[0,1]\times[0,1]\longrightarrow[0,1]$ be the function defined by
$G(x,y)=\left\{\begin{array}[]{ll}\alpha_{1}y&\text{ if }0\leqslant x\leqslant\frac{1}{2},\\ \alpha_{2}y&\text{ if }\frac{1}{2}<x\leqslant 1.\end{array}\right$.
It is easy to see that $G$ is discontinuous on the set $\{\frac{1}{2}\}\times[0,1]$. Moreover, $G$ satisfies (H2) since $|G|_{\zeta}=0$ (see Equation (13)), for all $\zeta$. Thus, $G$ is a $\alpha_{3}$-contraction, where $\alpha_{3}=\max\{\alpha_{1},\alpha_{2}\}$. Since $L=1$ we have that $(\alpha_{3}L)^{\zeta}<1$, for all $\zeta$. Therefore, $F$ satisfies all hypothesis (f1), (f2), (f3), (H1), (H2) and $(\alpha_{3}L)^{\zeta}<1$.
Example 2.6.
(Discontinuous Maps) Let $F=(f_{\alpha},G)$ be the measurable map, where $f_{\alpha}$ is again from the previous example 2.4. Consider a real number $0\leq\alpha_{2}<1$ and $\zeta$-Hölder functions $h_{1}:[0,\dfrac{1}{2}]\longrightarrow[0,1]$, $h_{2}:[\dfrac{1}{2},1]\longrightarrow[0,1]$ such that $h_{1}(\dfrac{1}{2})\neq h_{2}(\dfrac{1}{2})$ and $0\leq h_{1},h_{2}<\alpha_{2}<1$. Let $G:[0,1]\times[0,1]\longrightarrow[0,1]$ be the function defined by
$G(x,y)=\left\{\begin{array}[]{ll}h_{1}(x)y&\text{ if }0\leqslant x\leqslant\frac{1}{2},\\ h_{2}(x)y&\text{ if }\frac{1}{2}<x\leqslant 1.\end{array}\right$.
It is easy to see that $G$ is discontinuous on the set $\{\frac{1}{2}\}\times[0,1]$. Moreover, $G$ satisfies (H2) since $|G|_{\zeta}\leq\max\{|h_{1}|_{\zeta},|h_{2}|_{\zeta}\}$. Thus, $G$ is a $\alpha_{2}$-contraction. Since $L=1$ we have that $(\alpha_{2}L)^{\zeta}<1$. Therefore, $F$ satisfies all hypothesis (f1), (f2), (f3), (H1), (H2) and $(\alpha_{2}L)^{\zeta}<1$.
@@
where $l\geq 2$ is an integer, $0<\alpha<1$ is a real number, and $o$ is a $C^{2}$ function on $S^{1}$. In [23], M. Tsujii proved the existence of an ergodic probability measure $\mu$ on $S^{1}\times\mathbb{R}$ such that Lebesgue almost every point is generic. That is,
@@
Let us consider the setting: $M=S^{1}$, K=[2,2]K=[-2,2], $f=lx$, $G(x,y)=\alpha y+o(x)$, $\varphi=-\log|f^{\prime}|$ and $\Phi=\varphi\circ\pi_{1}$. Since $f^{\prime}$ is constant, it follows that $\Phi\in\mathcal{P}$ and $\varphi\in\mathcal{PH}$, according to definitions LABEL:P and LABEL:PH. Therefore, there exists a unique equilibrium state $\mu_{0}\in S^{1}\cap S^{\infty}$ for the pair $(F,\Phi)$. Moreover, all Theorems A through J apply to $(F,\mu_{0})$. Additionally, since $\varphi=-\log|f^{\prime}|$, the associated conformal measure $\nu$ coincides with Lebesgue measure $m$ on $S^{1}$. This implies $\pi_{1*}\mu_{0}<<m$, and by Theorem A of [7], $\pi_{1*}\mu_{0}$ is ergodic.
We claim that $\mu_{0}$ coincides with the physical measures, $\mu$, constructed in [23]. Indeed, since $\mu$ is ergodic and absolutely continuous with respect to the Lebesgue measure on $S^{1}\times K$, its projection $\pi_{1*}\mu$ is also ergodic and absolutely continuous with respect to $m$ on $S^{1}$. By uniqueness, we must have $\pi_{1*}\mu=\pi_{1*}\mu_{0}$, implying that both $\mu$ and $\mu_{0}$ project onto the same measure. It means that, $\mu$ and $\mu_{0}$ projects onto the same measure. By Proposition 2.8, this implies $\mu=\mu_{0}$. The same argument applies even when $\mu$ is singular with respect to Lebesgue measure, provided that $\mu$ is physical.

### v2

@@ 1. Introduction
  • Generalized Potentials and Dynamics: We extend the statistical analysis to a significantly broader class of potentials $\overline{\mathscr{P}}_{\Sigma}$ and dynamical systems than those treated in [3]. Notably, our results encompass potentials with explicit dependence on the fiber direction, thereby eliminating the restrictive requirement of fiber-constancy;
@@
Organization of the paper
The remainder of this article is organized as follows. In Section 2, we formalize the setting for piecewise partially hyperbolic maps and discuss several illustrative examples. Section 3 is devoted to preliminary concepts, specifically the construction of the anisotropic spaces $\mathcal{H}^{\zeta}_{m}$, $\mathbf{L}^{\infty}_{m}$, and $\mathbf{S}^{\infty}$, along with the required metric tools. In Section 4, we develop the functional-analytic framework for the transfer operator. Section 5 focuses on the statistical properties of the system, where we prove the exponential decay of correlations and the Central Limit Theorem; precisely, we provide the proofs of Theorems B, C, and D. In Section 6, we introduce the class of perturbations under consideration, describe the functional-analytic perturbative framework, and establish our main stability results, concluding with the proof of Theorem 51. Finally, in Section 7, we extend our analysis to equilibrium states associated with potentials that are not constant along fibres, providing the proofs of Theorems F, G, H, and I.
@@ 2. Settings and Examples
Throughout this article, we consider skew–product maps $F:M\times K\longrightarrow M\times K$ of the form $F=(f,G)$, where $f:M\longrightarrow M$ and $G:M\times K\longrightarrow K$ are measurable maps specified below. We fix a probability measure $m$ on the compact metric space $(M,d_{1})$ satisfying the hypotheses listed in (P1), and an arbitrary probability measure $m_{2}$ on the Borel $\sigma$–algebra of the compact metric space $(K,d_{2})$. For convenience, we set $\Sigma:=M\times K$.
@@
Definition 2.2.
Let $\mathscr{P}_{M}$ be the set of Hölder potentials satisfying the following condition, which is open with respect to the Hölder norm:
@@
    There exists $\epsilon_{\varphi}>0$ sufficiently small such that
    $\sup\varphi-\inf\varphi<\epsilon_{\varphi},$ (2)
@@
We assume that the constants $\epsilon_{\varphi}$ and $L$ satisfy
$\exp(\epsilon_{\varphi})\,\frac{(\deg(f)-q)\sigma^{-\alpha}+qL^{\alpha}\bigl[1+(L-1)^{\alpha}\bigr]}{\deg(f)}<1.$ (4)
@@ 2.2. Spectral gap for $\mathcal{L}_{\varphi}$
      For every $\varphi\in\mathscr{P}_{M}$, there exists an equilibrium state $m(=m(\varphi))$, such that $m=h\nu$, where $0<\inf h\leq\sup h<+\infty$, (see (P1.2)) $h,\frac{1}{h}\in B_{s}$ and $\mathcal{L}_{\varphi}(h)=\lambda h$, $\lambda>0$. Moreover, $\nu$ is a conformal measure with Jacobian $\lambda e^{-\varphi}$;
@@
      There exist normed spaces of real valued functions (not necessarily Banach), closed under the standard multiplication, $(B_{s},|\cdot|_{s})\subset(B_{w},|\cdot|_{w})$ satisfying $|\cdot|_{w}\leq|\cdot|_{s}$ such that for every $\varphi\in\mathscr{P}_{M}$ the spectrum of the Ruelle-Perron-Frobenius of $f:M\longrightarrow M$, $\mathcal{L}_{\varphi}:B_{s}\longrightarrow B_{s}$, defined by
@@
Moreover, the following theorems 2.3 and 2.4 hold, where the proof can be found in [4].
@@
Note that, the probability $m$ is an equilibrium state for a potential $\varphi\in\mathscr{P}_{M}$. Thus $\mu_{0}$ depends on the function $\Phi=\varphi\circ\pi_{1}\in\mathscr{P}_{\Sigma}$ (see equation (24)).
2.4. Examples
In this section, we present several examples illustrating the applicability of the results obtained in this work. In some cases, namely, Examples 2.1, 2.2, 2.3, 2.4, and 2.5, we analyze either the base dynamics $f$ or the fibre dynamics $G$ separately, without explicitly constructing the full skew-product $(f,G)$. By this, we mean that any combination of such base and fibre dynamics gives rise to a skew-product to which all Theorems A through I apply. On the other hand, Example 2.6 considers a complete skew-product structure where Theorems A through I are directly applicable. Example 2.7 illustrates a continuous dynamics scenario characterized by uniform contraction along all fibers while preserving a fixed horizontal fiber. Consequently, Theorems A through I apply in this setting. Furthermore, this case is of particular interest as the construction of the non-trivial class $\mathcal{S}$ presented in Section 7.1 is directly applicable here.
@@ 2.4. Examples
We present a general strategy to construct examples by perturbing either the identity map or an expanding map near the identity.
Let $f_{0}\colon\mathbb{T}^{d}\to\mathbb{T}^{d}$ be an expanding map. Choose a finite covering $\mathcal{P}$ of $\mathbb{T}^{d}$ and an element $P_{1}\in\mathcal{P}$ containing a periodic point $p$ (possibly a fixed point). We then define a perturbation $f$ of $f_{0}$ inside $P_{1}$, using a pitchfork bifurcation, so that $p$ becomes a saddle point for $f$.
By construction, $f$ coincides with $f_{0}$ on $P_{1}^{c}$, where uniform expansion is preserved. The perturbation can be chosen so as to satisfy condition (f1), ensuring that $f$ is not excessively contracting inside $P_{1}$ and that it remains topologically mixing.
It is worth noting that a perturbation with these properties does not always exist. However, whenever such a perturbation can be carried out, condition (f3) is also satisfied. In this case, the invariant measure $m$ is absolutely continuous with respect to the Lebesgue measure, which is expanding, conformal, and positive on open sets. As a consequence, the system admits no periodic attractors.
@@
Assume, in the setting of the previous example, that $f_{0}$ is diagonalizable, with eigenvalues
$1<1+a<\lambda,$
associated respectively with the eigenvectors $e_{1}$ and $e_{2}$, and that $x_{0}$ is a fixed point of $f_{0}$. Fix parameters $a,\varepsilon>0$ satisfying $\log\!\left(\frac{1+a}{1-a}\right)<\varepsilon$ and
$\operatorname{e}^{\varepsilon}\!\left(\frac{(\deg f_{0}-1)(1+a)^{-\zeta}+(1/(1-a))^{\zeta}\bigl[1+(a/(1-a))^{\zeta}\bigr]}{\deg f_{0}}\right)<1.$
Note that any smaller choice of $a>0$ still satisfies these inequalities.
Let $\mathcal{U}$ be a finite covering of $M$ by open domains of injectivity for $f_{0}$. By refining the covering if necessary, we may assume that $x_{0}=(m_{0},n_{0})$ belongs to a unique element $U\in\mathcal{U}$. Choose $r>0$ sufficiently small so that $B_{2r}(x_{0})\subset U$.
Define $\rho=\eta_{r}\ast g$, where $\eta_{r}(z)=r^{-2}\eta(z/r)$, $\eta$ is the standard mollifier, and
$g(m,n)=\begin{cases}\lambda(1-a),&\text{if }(m,n)\in B_{r}(x_{0}),\\[5.69054pt] \lambda(1+a),&\text{otherwise}.\end{cases}$
@@
$f(m,n)=\bigl(m_{0}+\lambda(m-m_{0}),\;n_{0}+(\rho(m,n)/\lambda)(n-n_{0})\bigr).$
With this construction, $x_{0}$ is a saddle point of $f$. Moreover, the required assumptions are satisfied with
$\mathcal{A}=B_{2r}(x_{0}),\qquad L_{1}=\frac{1}{1-a},\qquad\sigma=1+2a.$
The only non-trivial condition to verify is (f3).
To this end, observe that
@@
where
$S=\{z\in\mathbb{R}^{2}:\ x-z\in B_{r}(x_{0}),\ y-z\notin B_{r}(x_{0})\},$
and
$S^{\prime}=\{z\in\mathbb{R}^{2}:\ y-z\in B_{r}(x_{0}),\ x-z\notin B_{r}(x_{0})\}.$
Let $x,y\in\mathbb{R}^{2}$ and write |xy|=qr|x-y|=qr. Define the annulus
$A_{q}=\{z\in\mathbb{R}^{2}:\ 1-q<|z|<1\}.$
Then
@@
Since
$N:=\sup_{q>0}\frac{\eta(A_{q})}{q^{\zeta}}<\infty,$
we may choose $a>0$ sufficiently small so that
$\frac{2aN}{1-a}<\varepsilon.$
Consequently, $H_{\zeta}(\rho)<\varepsilon\,\inf\rho$, and condition (f3) follows.
Example 2.4 (Manneville–Pomeau map).
Let $\alpha\in(0,1)$ and consider the $C^{1+\alpha}$ local diffeomorphism $f_{\alpha}\colon[0,1]\to[0,1]$ defined by
$f_{\alpha}(x)=\begin{cases}x\bigl(1+2^{\alpha}x^{\alpha}\bigr),&\text{if }0\leq x\leq\tfrac{1}{2},\\[5.69054pt] 2x-1,&\text{if }\tfrac{1}{2}<x\leq 1.\end{cases}$
Conditions (f1) and (f2) are clearly satisfied. Moreover, by explicitly computing the two inverse branches of $f_{\alpha}$, one checks that $L=1$.
To verify condition (f3), consider the family of potentials $\{\varphi_{\alpha,t}\}_{t\in(-t_{0},t_{0})}$, defined by
$\varphi_{\alpha,t}=-t\log|Df_{\alpha}|,$
for $t_{0}>0$ sufficiently small. This defines a family of $C^{\alpha}$ potentials such that
$\{\varphi_{\alpha,t}\}_{t\in(-t_{0},t_{0})}\subset\mathscr{P}_{M}\quad\text{for each }\alpha\in(0,1),$
as we now verify.
Indeed, for any $x,y\in[0,1]$,
$|\varphi_{\alpha,t}(x)-\varphi_{\alpha,t}(y)|$ $=$ $|t|\bigl|\log|Df_{\alpha}(x)|-\log|Df_{\alpha}(y)|\bigr|$
$=$ $|t|\left|\log\frac{|Df_{\alpha}(x)|}{|Df_{\alpha}(y)|}\right|.$
Since $|Df_{\alpha}|$ is bounded away from zero and infinity on $(0,1]$ and has at most polynomial growth near $x=0$, we obtain the uniform bound
$\left|\log\frac{|Df_{\alpha}(x)|}{|Df_{\alpha}(y)|}\right|\leq\log(2+\alpha),$
which yields
$|\varphi_{\alpha,t}(x)-\varphi_{\alpha,t}(y)|\leq|t|\log(2+\alpha).$
Consequently, for $|t|$ small enough, the Hölder constant of $\varphi_{\alpha,t}$ satisfies the required bound in (f3), completing the verification.
Example 2.5 (Discontinuous maps).
Fix real numbers $\alpha_{1}$ and $\alpha_{2}$ such that
$0\leq\alpha_{1}<\alpha_{2}<1.$
Define $G\colon[0,1]\times[0,1]\to[0,1]$ by
$G(x,y)=\begin{cases}\alpha_{1}y,&\text{if }0\leq x\leq\tfrac{1}{2},\\[5.69054pt] \alpha_{2}y,&\text{if }\tfrac{1}{2}<x\leq 1.\end{cases}$
Clearly, $G$ is discontinuous along the set $\{\tfrac{1}{2}\}\times[0,1]$. Nevertheless, $G$ satisfies condition (H2), since for every $\zeta>0$ we have
$|G|_{\zeta}=0$
(see Equation (13)), as $G$ does not depend on the base variable inside each atom of the partition. Moreover, $G$ is uniformly contracting along the fibres, with contraction rate $\alpha_{3}=\max\{\alpha_{1},\alpha_{2}\}.$
@@
where $l\geq 2$ is an integer, $0<\alpha<1$ is a real number, and $o$ is a $C^{2}$ function on the unit circle $S^{1}$. In [23], M. Tsujii proved the existence of an ergodic probability measure $\mu$ on $S^{1}\times\mathbb{R}$ such that Lebesgue almost every point is generic. That is,
@@
Let us consider the setting: $M=S^{1}$ (unit circle), K=[2,2]K=[-2,2], $f=lx$, $G(x,y)=\alpha y+o(x)$, $\varphi=-\log|f^{\prime}|$ and $\Phi=\varphi\circ\pi_{1}$. Since $f^{\prime}$ is constant, it follows that $\Phi\in\mathscr{P}_{\Sigma}$ and $\varphi\in\mathscr{P}_{M}$, according to Definitions 4.1 and 2.2. Therefore, there exists a unique equilibrium state $\mu_{0}\in\mathbf{S}^{\infty}$ for the pair $(F,\Phi)$. Moreover, all Theorems A through D apply to $(F,\mu_{0})$. Additionally, since $\varphi=-\log|f^{\prime}|$, the associated conformal measure $\nu$ coincides with Lebesgue measure $m$ on $S^{1}$. This implies $\pi_{1*}\mu_{0}\ll m$, and by Theorem A of [7], $\pi_{1*}\mu_{0}$ is ergodic. Moreover, the measure $\mu_{0}$, whose existence is guaranteed by Theorem 4.16, coincides with the physical measure $\mu$ constructed in [23] (refer to [4] for further details).
The next example was introduced in [15], where it was shown that the non-wandering set of $F$ is partially hyperbolic for certain fixed parameters satisfying the conditions stated below. This class of maps was subsequently studied in [16], [18], [17], and [19]. In [18], the inverse map $F^{-1}$ was considered and shown to admit a skew product structure, where the base dynamics is strongly topologically mixing and non-uniformly expanding, while the fibre dynamics is uniformly contracting. In [19], the author established the existence of equilibrium states and proved stability results for this class of systems.
Example 2.7.
(Partially hyperbolic horseshoes) Consider the cube $R=[0,1]\times[0,1]\times[0,1]\subset\mathbb{R}^{3}$ and the parallelepipeds
Consider a map defined for $(x,y,z)\in R_{0}$ as
$F_{0}(x,y,z)=(\rho x,f(y),\beta z),$
where $0<\rho<{1/3}$, $\beta>6$ and
$f(y)=\frac{1}{1-\left(1-\frac{1}{y}\right)e^{-1}}.$
Consider also a map defined for $(x,y,z)\in R_{1}$ as
where $0<\sigma<{1/3}$ and $3<\beta_{1}<4$. Define the horseshoe map $F$ on $R$ as
$F|_{R_{0}}=F_{0},\quad F|_{R_{1}}=F_{1},$